OpenStudy (anonymous):

John and Tim are looking at the equation the square root of the quantity of 3 times x plus 4 equals square root of x . John says that the solution is extraneous. Tim says the solution is non-extraneous. Is John correct? Is Tim correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (ybarrap):

The only way to know is to check it $$ \large{ \sqrt{3x+4}=x\\ 3x+4=x^2\\ x^2-3x-4=0\\ (x-4)(x+1)=0 } $$ Answer is x=4 and x=-1. But how can the square root of something (1st equation) equal to -1 (assuming we are dealing with real numbers, not complex). Plugging in -1 into the original equation also does not work. So the only solution is x=4, x=-1 is extraneous. The problem is a dilemma because in the second step we squared both sides, introducing -1 as a possible solution, where initially it was not.

OpenStudy (ybarrap):

So in a sense they are both correct.